# AQM: A Refresh of the Abstract Qubit Model for Quantum Co-design

### v1

@@ V Conclusion
VI Acknowledgement
This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage (C2QA) under contract number DE-SC0012704, (Basic Energy Sciences, PNNL FWP 76274). The Pacific Northwest National Laboratory is operated by Battelle for the U.S. Department of Energy under Contract DE-AC05-76RL01830.
@@ Appendix A Supporting the AQM by trapped ions
The trapped ion systems support well on the AQM. For a thorough review of trapped ion systems for QC and QIP, we suggest referring to Refs. Häffner et al. (2008); Monroe and Kim (2013); Bruzewicz et al. (2019); Brown et al. (2021).
Qubit states.
@@
In a trapped ion system, each ion inside the trap can be used as a qubit. The quantum information can either be encoded into the hyperfine levels Harty et al. (2014); Gaebler et al. (2016); Ballance et al. (2016); Wang et al. (2017, 2021) or Zeeman sublevels of a same orbital Keselman et al. (2011); Ruster et al. (2016); Hilder et al. (2022), or other quantum states in the specific ion level structures Bermudez et al. (2017); Toyoda et al. (2010); Pogorelov et al. (2021).
Universal control.
@@
Depending on the type of ion qubits, the quantum manipulation schemes for qubit universal control of the ion qubits also vary. Specifically, for the hyperfine ion qubits, single-qubit gates can be implemented using either optical Raman transitions Hilder et al. (2022) or microwave drives Ospelkaus et al. (2011); Harty et al. (2014). The microwave Pauli-X and Pauli-Y rotations with fidelity $\sim 99.9999\%$ have been reported Harty et al. (2014). The microwave gates duration has been improved to $\sim 1~{}\mu$s without sacrificing much gate fidelity Leu et al. (2023). Raman transition-based single-qubit gates achieve fidelity $99.993\%$ in $7.5~{}\mu$Ballance et al. (2016). Using ultra-fast laser pulses to strongly drive Raman transition can achieve a $\pi$-pulse in $\sim 50$ ps with fidelity $\sim 99\%$ Campbell et al. (2010). These gates provide the necessary tools for universal control of single qubits.
@@
The coupling between two ions can be induced using the Coulomb interactions between them. There are several schemes to perform two-qubit gates between trapped ions, e.g., the Cirac-Zoller gate Cirac and Zoller (1995), the Mølmer-Sørensen (MS) gate Sørensen and Mølmer (1999), and the Leibfried geometric phase gate Leibfried et al. (2003). Specifically, MS gate utilizes the phonon modes of the trapped ion chain to mediate the coupling between different trapped ions, which is widely used in trapped ion systems Gaebler et al. (2016); Blümel et al. (2021); Moses et al. (2023); Jeon et al. (2023); Saner et al. (2023), while the fidelity of two-qubit gates can reach $99.9\%$ Ballance et al. (2016). The two-qubit gates can be performed in $10$ to $500~{}\mu$Gaebler et al. (2016); Ballance et al. (2016); Saner et al. (2023). The two-qubit entangling gates combined with the single-qubit universal control provide universal gate sets for QC.
Long coherence time.
@@
Depending on the type of encoding, the coherence time of the qubits can be different. For instance, in the Zeeman qubits, the coherence time can reach $300$ ms Ruster et al. (2016), while the hyperfine states are more coherent, and hence the coherence time of a hyperfine qubit can reach several minutes to an hour Harty et al. (2014); Wang et al. (2017, 2021) ($5500$ s reported in Ref. Wang et al. (2021)).
Qubit measurements.
@@
The qubit measurement is achieved using optical approaches, where one of the ion qubit states is excited to an optical active state, such that the relaxation of the ion can emit optical photons for detection. Therefore, by distinguishing the collected emitted photons, the state of the ion can be determined. This physical process is a projective measurement of the qubit computation states. The measurements on ion states can be obtained with fidelity $>99.9\%$ Burrell et al. (2010); Christensen et al. (2020); Todaro et al. (2021).
Connectivity.
@@
The ions are trapped using radio-frequency Paul traps Neuhauser et al. (1980) and other types of electromagnetic traps Dehmelt (1990); Chiaverini et al. (2005); Seidelin et al. (2006); Labaziewicz et al. (2008). The trapped ion systems can have all-to-all connectivity between the ions trapped in a single trap. This is due to the mechanism of two-qubit gate operations. The Mølmer-Sørensen gate requires using the collective motion (phonon) modes to mediate the coupling, which enables this all-to-all connectivity. However, limited by the size of the ion trap and the distinguishability between different phonon modes, it is impossible to have millions of ions trapped in a single trap and have all-to-all connectivity. Furthermore, due to the 1D nature of the ion traps, it is hard to maintain a higher-dimensional ion array to improve integration and connectivity.
@@
With the techniques of ion shuttling, it is possible to select the ion qubits that are necessary to be coupled in the quantum algorithm, and shuttling the ion qubits into the same ion trap to enable the coupling Moses et al. (2023). This technique potentially increases the connectivity of the ion qubits. However, shuttling operations can increase the decoherence noise and time consumption.
@@
In summary, trapped ion systems also support the AQM well. Similar to the superconducting qubit systems, the imperfection of the gate operations and decoherence limit the support to an ideal AQM. With the quantum error correction codes, it is possible to overcome the errors in the quantum operations. Compared with superconducting qubits, trapped ions can have longer coherence time, and all-to-all connectivity between ions in the same trap, which enables potential application to quantum memory and quantum simulation Zhang et al. (2017); Lv et al. (2018); Monroe et al. (2021); Kiesenhofer et al. (2023).
@@ Appendix B Supporting the AQM by neutral atoms
The trapped neutral atom system is becoming increasingly popular with the new progress of showing the potential of enabling error correction functionalities. Neutral atoms can also have good support for the AQM.
Qubit states.
@@
Neutral atom systems use the spin-electronic states of the trapped atoms to encode quantum information, similar to the ion qubits in the trapped ion system. There are a few strategies to encode quantum information into states of Rydberg atoms. For instance, the ground state of the atom and its Rydberg excited state can be used to encode the qubit state |$\ket{0}$ and |$\ket{1}$ states Levine et al. (2018); Madjarov et al. (2020), which is referred to as GR qubits in Ref. Picken et al. (2018); Morgado and Whitlock (2021), while two hyperfine ground states of the Rydberg atoms can also be used Sheng et al. (2018); Graham et al. (2019); Levine et al. (2019); Bluvstein et al. (2022), which is called GG qubits in Ref. Morgado and Whitlock (2021)). The Hilbert space spanned by the states in different selections can also be fully addressed.
Universal control.
@@
The single-qubit universal control can be achieved using optical Raman transitions or using microwave drives with optically activated Stack shifts Yavuz et al. (2006); Xia et al. (2015); Bluvstein et al. (2022); Levine et al. (2022). The Rabi rate of $2$ MHz has been realized for single-qubit gate operations Levine et al. (2022). With these control methods, the Hilbert space can be fully addressed.
@@
To support a universal gate set for QC, entangling gates between two qubits are necessary. The Rydberg interactions can be leveraged to perform entangling gates. When the atom is excited to a highly excited state (Rydberg state), the radius of the Rydberg state is much larger than the radius of the atom in the ground state, which activates a strong dipole-dipole interaction between Rydberg atoms. Specifically, the dipole interaction makes an excited Rydberg atom strongly shift the energy level of the other Rydberg atoms, which blocks the excitation of the nearby atoms. This effect is named “Rydberg blockade” Lukin et al. (2001); Jaksch et al. (2000); Urban et al. (2009), which enables fast two-qubit gates. The two-qubit gates with fidelity $97.4\%$ Levine et al. (2019); Bluvstein et al. (2022) and $99.5\%$ Evered et al. (2023); Bluvstein et al. (2023) have been demonstrated in experiments. Specifically, the CZ gate between two atoms can be implemented in $\sim 200$ ns, which greatly suppresses the error from the decoherence of the Rydberg state Evered et al. (2023).
Long coherence time.
@@
The coherence time of the qubits varies according to the species of qubits and the trapped atoms, ranging from a few microseconds to a few seconds Norcia et al. (2019); Ebadi et al. (2021); Jenkins et al. (2022); Ma et al. (2022). The coherence time can be further enhanced using dynamical decoupling to $1.5$ s for Rb atoms Bluvstein et al. (2022, 2023) and $3.7$ s for Yb atoms Jenkins et al. (2022).
Qubit measurements.
@@
The measurement of the trapped neutral atoms is similar to the method used in trapped ion systems. Fast measurements with descent measurement fidelity on the qubit state have been demonstrated, which can be used in quantum error correction Xu et al. (2021); Singh et al. (2023); Deist et al. (2022); Graham et al. (2023).
Connectivity.
@@
Compared to the trapped ion systems, the neutral atoms cannot be trapped using RF electromagnetic traps. Instead, neutral atoms are optically trapped into optical lattices. With the optical trapping techniques, higher dimensional optical lattices have been realized Wang et al. (2015, 2016); Xia et al. (2015); Maller et al. (2015); Graham et al. (2019). In addition, due to the large radius of the atom Rydberg state (can reach a few micrometers Urban et al. (2009); Gaëtan et al. (2009)), the entangling gates can be applied to two nonadjacent atoms, which potentially increase the connectivity of the neutral atom qubits.
@@
Furthermore, atoms can be precisely transported with high fidelity by adjusting the optical traps Ebadi et al. (2021); Bluvstein et al. (2024). Through this shuttling process, atoms trapped in distant locations can be brought to nearby sites, implementing entangling gate operations. This significantly enhances the connectivity of neutral atom systems.
@@
In essence, neutral atom systems show promise as a robust platform for AQM. Like superconducting qubits and trapped ions, a key challenge lies in developing error correction codes and demonstrating error suppression for full AQM support. However, the high connectivity of atoms in neutral atom systems makes them particularly suitable for quantum simulation applications Ebadi et al. (2021); Wu et al. (2021); Daley et al. (2022); Gross and Bloch (2017).

### v2

@@ 5 Conclusion
6 Acknowledgement
This work was solely supported by U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Codesign Center for Quantum Advantage (C2QA) under contract number DE-SC0012704, (Basic Energy Sciences, PNNL FWP 76274)
@@ Appendix A Supporting the AQM by trapped ions
Trapped ion systems support well on the AQM. For a thorough review of trapped ion systems for QC and QIP, we suggest referring to Refs. [75, 76, 77, 78].
a. Qubit states.
@@
In a trapped ion system, each ion inside the trap can be used as a qubit. The quantum information can either be encoded into the hyperfine levels [91, 92, 93, 94, 95] or Zeeman sublevels of a same orbital [249, 250, 112], or other quantum states in the specific ion level structures [85, 251, 86].
b. Universal control.
@@
Depending on the type of ion qubits, quantum manipulation schemes for qubit universal control of the ion qubits also vary. Specifically, for hyperfine ion qubits, single-qubit gates can be implemented using either optical Raman transitions [112] or microwave drives [113, 91]. The microwave Pauli-X and Pauli-Y rotations with fidelity $\sim 99.9999\%$ have been reported [91]. The microwave gates duration has been improved to $\sim 1\leavevmode\nobreak\ \mu$s without sacrificing much gate fidelity [252]. Raman transition-based single-qubit gates achieve fidelity $99.993\%$ in $7.5\leavevmode\nobreak\ \mu$[93]. Using ultra-fast laser pulses to strongly drive Raman transition can achieve a $\pi$-pulse in $\sim 50$ ps with fidelity $\sim 99\%$ [253]. These gates provide the necessary tools for universal control of single qubits.
@@
The coupling between two ions can be induced using the Coulomb interactions between them. There are several schemes to perform two-qubit gates between trapped ions, e.g., the Cirac-Zoller gate [254], the Mølmer-Sørensen (MS) gate [123], and the Leibfried geometric phase gate [255]. Specifically, MS gate utilizes the phonon modes of the trapped ion chain to mediate the coupling between different trapped ions, which is widely used in trapped ion systems [92, 256, 257, 258, 117], while the fidelity of two-qubit gates can reach $99.9\%$ [93]. The two-qubit gates can be performed in $10$ to $500\leavevmode\nobreak\ \mu$[92, 93, 117]. The two-qubit entangling gates combined with the single-qubit universal control provide universal gate sets for QC.
c. Long coherence time.
@@
Depending on the type of encoding, the coherence time of the qubits can be different. For instance, in the Zeeman qubits, the coherence time can reach $300$ ms [250], while hyperfine states are more coherent, and hence the coherence time of a hyperfine qubit can reach several minutes to an hour [91, 94, 95] ($5500$ s reported in Ref. [95]).
d. Qubit measurements.
@@
The qubit measurement is achieved using optical approaches, where one of the ion qubit states is excited to an optical active state, such that the relaxation of the ion can emit optical photons for detection. Therefore, by distinguishing the collected emitted photons, the state of the ion can be determined. This physical process is a projective measurement of the qubit computation states. The measurements on ion states can be obtained with fidelity $>99.9\%$ [259, 260, 261].
e. Connectivity.
@@
The ions are trapped using radio-frequency Paul traps [262] and other types of electromagnetic traps [263, 264, 265, 266]. The trapped ion systems can have all-to-all connectivity between the ions trapped in a single trap. This is due to the mechanism of two-qubit gate operations. The Mølmer-Sørensen gate requires using the collective motion (phonon) modes to mediate the coupling, which enables this all-to-all connectivity. However, limited by the size of the ion trap and the distinguishability between different phonon modes, it is impossible to have millions of ions trapped in a single trap and have all-to-all connectivity. Furthermore, due to the 1D nature of the ion traps, it is hard to maintain a higher-dimensional ion array to improve integration and connectivity.
@@
With the techniques of ion shuttling, it is possible to select the ion qubits that are necessary to be coupled in the quantum algorithm, and shuttling the ion qubits into the same ion trap to enable the coupling [257]. This technique potentially increases the connectivity of the ion qubits. However, shuttling operations can increase the decoherence noise and time consumption.
@@
In summary, trapped ion systems also support the AQM well. Similar to the superconducting qubit systems, the imperfection of the gate operations and decoherence limit the support to an ideal AQM. With the quantum error correction codes, it is possible to overcome the errors in the quantum operations. Compared with superconducting qubits, trapped ions can have longer coherence time, and all-to-all connectivity between ions in the same trap, which enables potential application to quantum memory and quantum simulation [267, 66, 65, 268].
@@ Appendix B Supporting the AQM by neutral atoms
Trapped neutral atom systems are becoming increasingly popular with the new progress of showing the potential of enabling error correction functionalities. Neutral atoms can also have good support for the AQM.
a. Qubit states.
@@
Neutral atom systems use the spin-electronic states of the trapped atoms to encode quantum information, similar to the ion qubits in the trapped ion system. There are a few strategies to encode quantum information into states of Rydberg atoms. For instance, the ground state of the atom and its Rydberg excited state can be used to encode the qubit state |$\ket{0}$ and |$\ket{1}$ states [87, 88], which is referred to as GR qubits in Ref. [89, 90], while two hyperfine ground states of the Rydberg atoms can also be used [96, 97, 98, 81], which is called GG qubits in Ref. [90]). The Hilbert space spanned by the states in different selections can also be fully addressed.
b. Universal control.
@@
The single-qubit universal control can be achieved using optical Raman transitions or using microwave drives with optically activated Stack shifts [114, 115, 81, 116]. The Rabi rate of $2$ MHz has been realized for single-qubit gate operations [116]. With these control methods, the Hilbert space can be fully addressed.
@@
To support a universal gate set for QC, entangling gates between two qubits are necessary. The Rydberg interactions can be leveraged to perform entangling gates. When the atom is excited to a highly excited state (Rydberg state), the radius of the Rydberg state is much larger than the radius of the atom in the ground state, which activates a strong dipole-dipole interaction between Rydberg atoms. Specifically, the dipole interaction makes an excited Rydberg atom strongly shift the energy level of the other Rydberg atoms, which blocks the excitation of the nearby atoms. This effect is named “Rydberg blockade” [269, 270, 271], which enables fast two-qubit gates. The two-qubit gates with fidelity $97.4\%$ [98, 81] and $99.5\%$ [118, 119] have been demonstrated in experiments. Specifically, the CZ gate between two atoms can be implemented in $\sim 200$ ns, which greatly suppresses the error from the decoherence of the Rydberg state [118].
c. Long coherence time.
@@
The coherence time of the qubits varies according to the species of qubits and the trapped atoms, ranging from a few microseconds to a few seconds [272, 67, 273, 274]. The coherence time can be further enhanced using dynamical decoupling to $1.5$ s for Rb atoms [81, 119] and $3.7$ s for Yb atoms [273].
d. Qubit measurements.
@@
The measurement of the trapped neutral atoms is similar to the method used in trapped ion systems. Fast measurements with descent measurement fidelity on the qubit state have been demonstrated, which can be used in quantum error correction [275, 276, 277, 278].
e. Connectivity.
@@
Compared to trapped ion systems, the neutral atoms cannot be trapped using RF electromagnetic traps. Instead, neutral atoms are optically trapped into optical lattices. With the optical trapping techniques, higher dimensional optical lattices have been realized [279, 280, 115, 281, 97]. In addition, due to the large radius of the atom Rydberg state (can reach a few micrometers [271, 282]), the entangling gates can be applied to two nonadjacent atoms, which potentially increase the connectivity of the neutral atom qubits.
@@
Furthermore, atoms can be precisely transported with high fidelity by adjusting the optical traps [67, 12]. Through this shuttling process, atoms trapped in distant locations can be brought to nearby sites, implementing entangling gate operations. This significantly enhances the connectivity of neutral atom systems.
@@
In essence, neutral atom systems show promise as a robust platform for AQM. Like superconducting qubits and trapped ions, a key challenge lies in developing error correction codes and demonstrating error suppression for full AQM support. However, the high connectivity of atoms in neutral atom systems makes them particularly suitable for quantum simulation applications [67, 79, 68, 283].